# Performant Automatic BLAS Offloading on Unified Memory Architecture with OpenMP First-Touch Style Data Movement

### v2

@@ 4.4.2. Impact of page sizes on counter-based migration
Here I again run simple dgemm tests to understand the performance and issues. Notice that the aforementioned counter-based page migration mechanism does not work with the 4KB page size, so one can measure the performance of GPU kernel running on LPDDR5X, while this cannot be done when page size is 64KB as matrices are partially migrated to GPU as explained in Section 4.4.1.
The test results are summarized in Table 6. It can be seen that CPU accessing HBM3 memory is substantially slower under 64KB page than 4KB page for both problem sizes. There is a substantial difference even for CPU accessing LPDDR5 for the second workload, runtime with 64KB page is 15.8ms, much slower than the 10.9ms under 4KB page.
@@ 4.4.3. Sensitivity of page alignment for GPU accessing system allocated HBM
As seen above in application tests, the BLAS runtime under Device First-Use policy is noticeably slower than the BLAS running on the cudaMalloc memory created in the Mem-Copy policy. Further investigation shows that the CUDA kernel is slow when running on system malloc allocated HBM, unless the matrices are aligned to the page. Table 7 shows the difference with page size being 64 KB. All matrices are allocated by malloc, and pinned to HBM by using numactl -m 1 ./exe. When the matrices are aligned to page size, cublasDgemm speed can be up to nearly 50% faster than if data is not aligned. The impact is more significant in memory bandwidth bound kernels. The aligned performance is identical to that of the same CUDA kernel executing on cudaMalloc memory.
Application tests reveal that BLAS runtime performance under the Device First-Use policy is noticeably slower compared to BLAS operating on cudaMalloc memory used in the Mem-Copy policy. Further investigation indicates that CUDA kernel performance is suboptimal when operating on system-allocated HBM3 via malloc unless the matrices are aligned to the page. Table 7 highlights the performance differences. The tests use 64KB page size. All matrices are allocated using malloc and pinned to HBM3 with the command numactl -m 1 ./exe. When matrices are page-aligned, the performance of cublasDgemm improves by nearly 50% compared to cases where the data is not aligned. This improvement is particularly pronounced for memory bandwidth-bound kernels. In the page-aligned case, the performance on HBM3 allocated by malloc matches that of the same CUDA kernel executing on cudaMalloc memory. The reason of such behavior is unknown, and it partially defeats the advantage of unified memory architecture.
@@ 5. Conclusion and Future Work
In this paper, a previous proof-of-concept BLAS auto-offload prototype tool is further optimized and extended to all level-3 BLAS operations. To overcome the data transfer bottleneck between CPU and GPU, an OpenMP first-touch type of data management strategy, namely GPU first-use policy, is proposed and discussed in details. The new policy minimizes the data transfer between CPU and GPU in practical BLAS-heavy scientific computing codes, such as the quantum chemistry and quantum physics applications. Performance testings for several such codes show production level GPU performance. As all the performance evaluations were carried out on the NVIDIA Grace-Hopper architecture, limitations and advantages of the memory subsystem in Grace-Hopper are outlined. The proposed GPU first-use data management policy is applicable not limited to the architecture tested here, but can be universally applied for any CPU-GPU architecture that allows cache-coherent access of the GPU memory from CPU. This tool enables domain scientists to rapidly adopt modern GPUs in complex or legacy, BLAS-heavy codes to speedup scientific discoveries, and can also be used to quickly assess the potential benefits of GPU acceleration before a formal effort of porting. Work is in progress is to support auto BLAS offload on AMD GPUs. Additionally, FFTW auto-offload is under exploration.

### v3

@@ 4.4.2. Impact of page sizes on counter-based migration
Here we again run simple dgemm tests to understand the performance and issues. Notice that the aforementioned counter-based page migration mechanism does not work with the 4KB page size, so one can measure the performance of GPU kernel running on LPDDR5X, while this cannot be done when page size is 64KB as matrices are partially migrated to GPU as explained in Section 4.4.1.
The test results are summarized in Table 7. It can be seen that CPU accessing HBM3 memory is substantially slower under 64KB page than 4KB page for both problem sizes. There is a substantial difference even for CPU accessing LPDDR5X for the second workload, runtime with 64KB page is 15.8ms, much slower than the 10.9ms under 4KB page.
@@ 4.4.3. Sensitivity of page alignment for GPU accessing system allocated HBM
Application tests reveal that BLAS runtime performance under the Device First-Use policy is noticeably slower compared to BLAS operating on cudaMalloc memory used in the Mem-Copy policy. Further investigation indicates that CUDA kernel performance is suboptimal when operating on system-allocated HBM3 via malloc unless the matrices are aligned to the page. Table 8 highlights the performance differences. The tests use 64KB page size. All matrices are allocated using malloc and pinned to HBM3 with the command numactl -m 1 ./exe. When matrices are page-aligned, the performance of cublasDgemm improves by nearly 50% compared to cases where the data is not aligned. This improvement is particularly pronounced for memory-bandwidth-bound kernels. In the page-aligned case, the performance on HBM3 allocated by malloc matches that of the same CUDA kernel executing on cudaMalloc memory. The reason of such behavior is unknown, and it partially defeats the advantage of unified memory architecture.
@@ 5. Conclusion and Future Work
In this paper, a previous proof-of-concept BLAS auto-offload prototype tool is further optimized and extended to all level-3 BLAS operations. To overcome the data transfer bottleneck between CPU and GPU, an OpenMP first-touch type of data management strategy, namely the GPU first-use policy, is proposed and discussed in details. The new policy minimizes the data transfer between CPU and GPU in practical BLAS-heavy scientific computing codes, such as the quantum chemistry and quantum physics applications. Performance tests for several such codes show production-level GPU performance. As all performance evaluations were performed on the NVIDIA Grace-Hopper architecture, limitations and advantages of the memory subsystem in Grace-Hopper are outlined. The proposed GPU first-use data management policy is applicable not only to the architecture tested here but can be universally applied for any CPU-GPU architecture that allows cache-coherent access of the GPU memory from CPU. This tool enables domain scientists to rapidly adopt modern GPUs in complex or legacy, BLAS-heavy codes to speed up scientific discoveries, and can also be used to quickly assess the potential benefits of GPU acceleration before a formal effort of porting. The work in progress is to support auto BLAS offload on AMD GPUs. Additionally, FFTW auto-offload is under exploration.